6. An object is moving at a velocity of 30 m/s. It accelerates to a velocity of 5 m/s over a time of 12.5 s. What acceleration did it experience?

Formula: a = (v - v0)

---------

t

 5. A WW Beetle is traveling at a constant velocity of 12 mu's when it experience of 2.5 m/s for a time of 20 What was velocity be after that acceleration?

Formula: v = v0 + at


10. A tiger races after a deer with a constant acceleration of seconds was it running for? Formula x = 1/2 at2


11. What is the velocity of a dropped iron after it has fallen for 95?

Formula: v= v0 + gt






1
Expert's answer
2022-01-16T13:08:50-0500

6. The acceleration is


"a=\dfrac{30-5}{12.5}=2\text{ m/s}^2."


5. After the acceleration at 2.5 m/s2 during 20 s its final speed is


"v=12+2.5·20=62\text{ m/s}."


10. The time can be found from the following relation:


"t=\sqrt\dfrac {2x}a."


11. The final velocity can be found by the following relation:


"v=v_0+at."

To obtain the numerical answer, please substitute the values in the equation above.